# Deep learning for dynamic graphs: models and benchmarks

### 27 Dec 2023 v2

@@ II-B1 Spectral convolution
$\small\mathbf{x}^{\ell+1}_{u}=\mathbf{\Theta}_{0}\mathbf{x}_{u}^{\ell}+\mathbf{\Theta}_{1}\sum_{v\in\mathcal{N}_{u}}\frac{1}{\sqrt{\mathbf{deg}(v)\,\mathbf{deg}(u)}}\mathbf{x}^{\ell}_{v},$ (2)
@@ III Representation learning for dynamic Graphs: Notation and Taxonomy
A dynamic graph (also referred to as temporal graph) is a tuple $\mathcal{G}(t)=(\mathcal{V}(t),\mathcal{E}(t),\mathbf{X}(t),\mathbf{E}(t))$, defined for $t\geq 0$. Differently from static graphs, all elements in the tuple are functions of time $t$. Thus, $\mathcal{V}(t)$ provides the set of nodes which are present in the graph at time $t$, and $\mathcal{E}(t)\subseteq\{\{u,v\}\,|\,u,v\in\mathcal{V}(t)\}$ defines the links between them. Analogously, $\mathbf{X}(t)$ and $\mathbf{E}(t)$ define node states and edge attributes at time $t$. Although, $\mathcal{V}(t)$ can theoretically change over time, in practice it is often considered fixed for the ease of computation, which means that all the nodes that will appear in the dynamic graph are known in advance. Hence, $\mathcal{V}(t)=\mathcal{V}$ for $t\geq 0$.
@@
Generally, the temporal neighborhood of a node $u$ at time $t$, consists of all the historical neighbors of $u$, prior to current time $t$, i.e., $\mathcal{N}^{t}_{u}=\{v\in\mathcal{V}(t)\,|\,(t_{i},\,interaction,\,v,u)\in\mathcal{G}\wedge t_{i}<t\}$.
@@ V-2 Stacked architectures
In the case of sequential encoding of spatial and temporal information, Xu et al. [12] introduce TGAT [12], a model that learns the parameters of a continuous function that characterize the continuous-time stream. Similar to GraphSAGE and GAT models, TGAT employs a local aggregator that takes as input the temporal neighborhood and the timestamp and computes a time-aware embedding of the target node by exploiting an attention mechanism. The $\ell$-th layer of TGAT computes the temporal embedding of node $u$ at time $t$ as
@@ VI The benchmark problem
With this in mind, we designed three benchmarks to assess models that deal with spatio-temopral graphs, general D-TDGs, and C-TDGs. Specifically, we evaluated methods for D-TDGs on both link and node prediction tasks while we constrained the evaluation of C-TDG methods to link prediction tasks due to the scarcity of suitable datasets. To do so, we extended the library PyDGN [69] to the D-TDG learning setting to foster reproducibility and robustness of results. With the same aim, we developed a Pytorch Geometric [70] based framework to allow reproducible results in the continuous scenario. Lastly, in the Supplementary materials (Appendix A) we provide the community with a selection of datasets useful for benchmarking future works. An interest reader is referred to SNAP [71], TSL [72], and Network Repository [73] for a broader data collections.
@@ Results
Overall, DCRNN and GCRN-GRU achieve the better performance on the selected tasks. Interestingly, they both rely on Chebyshev spectral convolution and GRU, but with different architectural structure. Indeed, DCRNN employs a stacked architecture, while GCRN-GRU embeds the DGN into the RNN, enabling a combined modeling of the temporal and spatial information. This result shows that there is not a superior architectural design, in these tasks. However, it seems relevant to include a bigger neighborhood in the computation (e.g., by exploiting a larger Chebishev polynomial filter size). Indeed, even if A3TGCN employs an attention mechanism to capture more global information, it is not enough to achieve comparable performance to DCRNN or GCRN-based approaches. Nevertheless, it is noteworthy that the superior performance of these approaches comes at the expense of computational speed, as it is shown in the Supplementary materials (Appendix C).
@@
Table IV shows the results on general D-TDGs. Differently than the spatio-temporal setting, different tasks benefit from different architectures. Indeed, integrating topology’s changes (such as in GCLSTM and LRGCN) is more effective in link prediction tasks, while evolving the parameters of the DGN is more beneficial for node-level tasks, since it is more difficult to change the parameters of a static DGN to predict the topological evolution of the system. Notably, DynGESN achieves comparable results by exploiting only few trainable parameters and reduced computational overhead (see Supplementary materials, Appendix C), showing an advantageous tradeoff between performance and complexity. This makes it an ideal choice when the computational resources are limited.
@@
MAE MSE Balanced Acc F1 AUROC F1 AUROC F1
@@
Overall, TGN generally outperforms all the other methods, showing consistent improvements over DyRep and JODIE. This result shows how the spatial information is fundamental for the effective resolution of the tasks. Indeed, an advantage of TGAT and TGN is that they can exploit bigger neighborhoods with respect to DyRep, which uses the information coming from one-hop distance, and JODIE, which only encode the source and destination information. Despite these results, we observe that the temporal information is still extremely relevant to achieve good performance. In fact, the EdgeBank baseline is able to exceed 91% AUROC score by only looking at the graph’s history. This is even more evident in the LastFM task, which, as observed in [80], contains more reoccurring edges with respect to Wikipedia and Reddit. Consequently, such a task is comparatively easier to solve by solely exploiting these temporal patterns. Considering that EdgeBank’s performance is directly correlated to the number of memorized edges, in this task, it is able to outperform all the other methods. Lastly, it is worth mentioning that the enhanced performance of TGN and TGAT is accompanied by a trade-off in computational speed, as illustrated in the Supplementary materials (Appendix C).
@@ Acknowledgements
Name Cit. Type Addition Deletion Addition Deletion
@@
embedding dim lr weight decay filter size normalization
@@
embedding dim lr weight decay n. DGN layers DGN dim
@@
AUROC F1 Balanced Acc AUROC F1 Balanced Acc
@@
AUROC F1 Acc AUROC F1 Acc AUROC F1 Acc
@@
Montevideo MetrLA PeMSBay Traffic min max min max min max min max A3TGCN ed: 0.009 lr: 0.015 wd: 0.099 ed: 0.127 ed: 0.039 lr: 0.081 ed: 0.006 lr: 0.009 DCRNN ed: 0.016 lr: 0.025 wd: 0.019 ed: 0.034 ed: 0.050 fs: 0.131 wd: 0.017 ed: 0.020 GCRN-GRU lr: 0.038 K: 0.043 wd: 0.057 K: 0.071 ed: 0.086 K: 0.150 wd: 0.040 K: 0.047 GCRN-LSTM wd: 0.031 ed: 0.039 wd: 0.056 ed: 0.082 ed: 0.072 K: 0.155 wd: 0.024 lr: 0.031 TGCN ed: 0.013 lr: 0.018 wd: 0.099 ed: 0.120 ed: 0.041 lr: 0.087 ed: 0.006 lr: 0.009
@@
Twitter tennis Elliptic AS-733 Bitcoin-$\alpha$ min max min max min max min max DynGESN lr: 0.004 $\sigma$: 0.006 lr: 0.004 ed: 0.005 ed: 0.003 $\sigma$: 0.005 $\gamma$: 0.009 lr: 0.052 EvolveGCN-H wd: 0.011 lr: 0.012 lr: 0.003 ed: 0.005 wd: 0.082 ed: 0.109 ns: 0.018 ed: 0.020 EvolveGCN-O ed: 0.010 lr: 0.012 lr: 0.005 ed: 0.007 ed: 0.095 wd: 0.105 lr: 0.022 wd: 0.024 GCLSTM lr: 0.012 ed: 0.016 wd: 0.004 ed: 0.005 K: 0.201 ed: 0.226 K: 0.025 lr: 0.207 LRGCN nb: 0.008 ed: 0.01 wd: 0.005 nb: 0.005 nb: 0.036 lr: 0.057 nb: 0.003 ed: 0.004
@@
Wikipedia Reddit LastFM min max min max min max DyRep ed: 0.004 wd: 0.009 lr: 0.002 wd: 0.003 lr: 0.004 ed: 0.007 JODIE lr: 0.023 ed: 0.027 wd: 0.007 ed: 0.007 wd: 0.004 ed: 0.006 TGAT gl: 0.004 re: 0.023 ge: 0.002 re: 0.011 gl: 0.007 wd: 0.030 TGN ed: 0.004 wd: 0.005 wd: 0.004 ge: 0.006 gl: 0.012 lr: 0.033
@@
Montevideo Metr-LA PeMSBay Traffic
@@
Twitter tennis Elliptic AS-773 Bitcoin $\alpha$
@@
Wikipedia Reddit LastFM

### 18 Mar 2024 v3

@@ II-B1 Spectral convolution
$\small\mathbf{x}^{\ell+1}_{u}=\sigma\left(\mathbf{\Theta}_{0}\mathbf{x}_{u}^{\ell}+\mathbf{\Theta}_{1}\sum_{v\in\mathcal{N}_{u}}\frac{\mathbf{x}^{\ell}_{v}}{\sqrt{\mathbf{deg}(v)\,\mathbf{deg}(u)}}\right),$ (2)
@@ III Representation learning for dynamic Graphs: Notation and Taxonomy
A dynamic graph (also referred to as temporal graph) is a tuple $\mathcal{G}(t)=(\mathcal{V}(t),\mathcal{E}(t),\mathbf{X}(t),\mathbf{E}(t))$, defined for $t\geq 0$. Differently from static graphs, all elements in the tuple are functions of time $t$. Thus, $\mathcal{V}(t)$ provides the set of nodes which are present in the graph at time $t$, and $\mathcal{E}(t)\subseteq\{\{u,v,t_{i}\}\,|\,u,v\in\mathcal{V}(t),t_{i}<t\}$ defines the links between them. Analogously, $\mathbf{X}(t)$ and $\mathbf{E}(t)$ define node states and edge attributes at time $t$. Although, $\mathcal{V}(t)$ can theoretically change over time, in practice it is often considered fixed for the ease of computation, which means that all the nodes that will appear in the dynamic graph are known in advance. Hence, $\mathcal{V}(t)=\mathcal{V}$ for $t\geq 0$.
@@
Generally, the temporal neighborhood of a node $u$ at time $t$, consists of all the historical neighbors of $u$, prior to current time $t$, i.e., $\mathcal{N}^{t}_{u}=\{(v,t_{i})\,|\,\{u,v,t_{i}\}\in\mathcal{E}(t),t_{i}<t\}$.
@@ V-2 Stacked architectures
In the case of sequential encoding of spatial and temporal information, Xu et al. [12] introduce TGAT, a model that learns the parameters of a continuous function that characterize the continuous-time stream. Similar to GraphSAGE and GAT models, TGAT employs a local aggregator that takes as input the temporal neighborhood and the timestamp and computes a time-aware embedding of the target node by exploiting an attention mechanism. The $\ell$-th layer of TGAT computes the temporal embedding of node $u$ at time $t$ as
@@ VI The benchmark problem
With this in mind, we designed three benchmarks to assess models that deal with spatio-temopral graphs, general D-TDGs, and C-TDGs. Specifically, we evaluated methods for D-TDGs on both link and node prediction tasks while we constrained the evaluation of C-TDG methods to link prediction tasks due to the scarcity of suitable datasets. To do so, we extended the library PyDGN [69] to the D-TDG learning setting to foster reproducibility and robustness of results. With the same aim, we developed a Pytorch Geometric [70] based framework to allow reproducible results in the continuous scenario. Lastly, in the Supplementary materials (Appendix -A) we provide the community with a selection of datasets useful for benchmarking future works. An interest reader is referred to SNAP [71], TSL [72], and Network Repository [73] for a broader data collections.
@@ Results
Overall, DCRNN and GCRN-GRU achieve the better performance on the selected tasks. Interestingly, they both rely on Chebyshev spectral convolution and GRU, but with different architectural structure. Indeed, DCRNN employs a stacked architecture, while GCRN-GRU embeds the DGN into the RNN, enabling a combined modeling of the temporal and spatial information. This result shows that there is not a superior architectural design, in these tasks. However, it seems relevant to include a bigger neighborhood in the computation (e.g., by exploiting a larger Chebishev polynomial filter size). Indeed, even if A3TGCN employs an attention mechanism to capture more global information, it is not enough to achieve comparable performance to DCRNN or GCRN-based approaches. Nevertheless, it is noteworthy that the superior performance of these approaches comes at the expense of computational speed, as it is shown in the Supplementary materials (Appendix -C).
@@
Table IV shows the results on general D-TDGs. Differently than the spatio-temporal setting, different tasks benefit from different architectures. Indeed, integrating topology’s changes (such as in GCLSTM and LRGCN) is more effective in link prediction tasks, while evolving the parameters of the DGN is more beneficial for node-level tasks, since it is more difficult to change the parameters of a static DGN to predict the topological evolution of the system. Notably, DynGESN achieves comparable results by exploiting only few trainable parameters and reduced computational overhead (see Supplementary materials, Appendix -C), showing an advantageous tradeoff between performance and complexity. This makes it an ideal choice when the computational resources are limited.
@@
Model MAE MSE Balanced Acc F1 AUROC F1 AUROC F1
@@
Overall, TGN generally outperforms all the other methods, showing consistent improvements over DyRep and JODIE. This result shows how the spatial information is fundamental for the effective resolution of the tasks. Indeed, an advantage of TGAT and TGN is that they can exploit bigger neighborhoods with respect to DyRep, which uses the information coming from one-hop distance, and JODIE, which only encode the source and destination information. Despite these results, we observe that the temporal information is still extremely relevant to achieve good performance. In fact, the EdgeBank baseline is able to exceed 91% AUROC score by only looking at the graph’s history. This is even more evident in the LastFM task, which, as observed in [80], contains more reoccurring edges with respect to Wikipedia and Reddit. Consequently, such a task is comparatively easier to solve by solely exploiting these temporal patterns. Considering that EdgeBank’s performance is directly correlated to the number of memorized edges, in this task, it is able to outperform all the other methods. Lastly, it is worth mentioning that the enhanced performance of TGN and TGAT is accompanied by a trade-off in computational speed, as illustrated in the Supplementary materials (Appendix -C).
@@ Acknowledgements
Model Cit. Type Addition Deletion Addition Deletion
@@
Model embedding dim lr weight decay filter size normalization
@@
Model embedding dim lr weight decay n. DGN layers DGN dim
@@
Model AUROC F1 Balanced Acc AUROC F1 Balanced Acc
@@
Model AUROC F1 Acc AUROC F1 Acc AUROC F1 Acc
@@
Montevideo MetrLA PeMSBay Traffic Model min max min max min max min max A3TGCN ed: 0.009 lr: 0.015 wd: 0.099 ed: 0.127 ed: 0.039 lr: 0.081 ed: 0.006 lr: 0.009 DCRNN ed: 0.016 lr: 0.025 wd: 0.019 ed: 0.034 ed: 0.050 fs: 0.131 wd: 0.017 ed: 0.020 GCRN-GRU lr: 0.038 K: 0.043 wd: 0.057 K: 0.071 ed: 0.086 K: 0.150 wd: 0.040 K: 0.047 GCRN-LSTM wd: 0.031 ed: 0.039 wd: 0.056 ed: 0.082 ed: 0.072 K: 0.155 wd: 0.024 lr: 0.031 TGCN ed: 0.013 lr: 0.018 wd: 0.099 ed: 0.120 ed: 0.041 lr: 0.087 ed: 0.006 lr: 0.009
@@
Twitter tennis Elliptic AS-733 Bitcoin-$\alpha$ Model min max min max min max min max DynGESN lr: 0.004 $\sigma$: 0.006 lr: 0.004 ed: 0.005 ed: 0.003 $\sigma$: 0.005 $\gamma$: 0.009 lr: 0.052 EvolveGCN-H wd: 0.011 lr: 0.012 lr: 0.003 ed: 0.005 wd: 0.082 ed: 0.109 ns: 0.018 ed: 0.020 EvolveGCN-O ed: 0.010 lr: 0.012 lr: 0.005 ed: 0.007 ed: 0.095 wd: 0.105 lr: 0.022 wd: 0.024 GCLSTM lr: 0.012 ed: 0.016 wd: 0.004 ed: 0.005 K: 0.201 ed: 0.226 K: 0.025 lr: 0.207 LRGCN nb: 0.008 ed: 0.01 wd: 0.005 nb: 0.005 nb: 0.036 lr: 0.057 nb: 0.003 ed: 0.004
@@
Wikipedia Reddit LastFM Model min max min max min max DyRep ed: 0.004 wd: 0.009 lr: 0.002 wd: 0.003 lr: 0.004 ed: 0.007 JODIE lr: 0.023 ed: 0.027 wd: 0.007 ed: 0.007 wd: 0.004 ed: 0.006 TGAT gl: 0.004 re: 0.023 ge: 0.002 re: 0.011 gl: 0.007 wd: 0.030 TGN ed: 0.004 wd: 0.005 wd: 0.004 ge: 0.006 gl: 0.012 lr: 0.033
@@
Model Montevideo Metr-LA PeMSBay Traffic
@@
Model Twitter tennis Elliptic AS-773 Bitcoin $\alpha$
@@
Model Wikipedia Reddit LastFM